OpenStudy (anonymous):

For the following true conditional statement, write the converse. If the converse is also true combine the statements as a biconditional. If x=3 then x+6=9 a.If x + 6 = 9, then x = 3. True; x = 3 if and only if x + 6 = 9 b.If x = 3 then x+6=9. False c.If x+6=9 then x=3, False d.If x+6=9 then then x=3. False; x+6=9 if and only if x=3

OpenStudy (anonymous):

I need help!

OpenStudy (anonymous):

if \( x=3\) then \(x+6=9 \)

OpenStudy (anonymous):

the converse, replace switch the order and get if \(x+6=9\) then \(x=3\)

OpenStudy (anonymous):

the converse it true as well that means \(x=3\) if and only if \(x=6=9\) usually abbreviated as \[x+3\iff x+6=9\]